### v2

# SegQuant: A Semantics-Aware and Generalizable Quantization Framework for Diffusion Models

Abstract
Diffusion models have demonstrated exceptional generative capabilities but are computationally intensive, posing significant challenges for deployment in resource-constrained or latency-sensitive environments. Quantization offers an effective means to reduce model size and computational cost, with post-training quantization (PTQ) being particularly appealing due to its compatibility with pre-trained models without requiring retraining or training data. However, existing PTQ methods for diffusion models often rely on architecture-specific heuristics that limit their generalizability and hinder integration with industrial deployment pipelines. To address these limitations, we propose SegQuant, a unified quantization framework that adaptively combines complementary techniques to enhance cross-model versatility. SegQuant consists of a segment-aware, graph-based quantization strategy (SegLinear) that captures structural semantics and spatial heterogeneity, along with a dual-scale quantization scheme (DualScale) that preserves polarity-asymmetric activations, which is crucial for maintaining visual fidelity in generated outputs. SegQuant is broadly applicable beyond Transformer-based diffusion models, achieving strong performance while ensuring seamless compatibility with mainstream deployment tools.
Introduction
Diffusion models (Ho, Jain, and Abbeel 2020) have emerged as a dominant class of generative models, demonstrating impressive performance across various applications including image synthesis, inpainting, video generation, etc. Despite their impressive performance, deploying diffusion models at scale remains challenging, particularly in high-concurrency settings where service providers must balance computational efficiency with output quality.
To address these challenges, quantization (Gholami et al. 2021) has emerged as a practical solution for reducing the computational burden of diffusion inference. One promising approach to improving deployment efficiency is post-training quantization (Jacob et al. 2018) (PTQ), which reduces model precision without requiring retraining or extensive fine-tuning. Quantization significantly improves inference speed and memory usage. However, reducing numerical precision often leads to performance degradation, especially in complex models like diffusion models that rely on iterative refinement over multiple denoising steps. Therefore, developing effective quantization strategies that preserve generation fidelity while maximizing efficiency has become a critical research direction.
Recent efforts in PTQ for diffusion models (Li et al. 2023; Huang et al. 2024; Chu et al. 2024; Wu et al. 2024) have focused on exploiting structural properties and temporal dynamics unique to these models. Meanwhile, mainstream industrial quantization frameworks (Paszke et al. 2019; Abadi et al. 2016; NVIDIA Corporation 2023) emphasize compatibility and modularity across diverse architectures, often at the expense of domain-specific optimizations. This divergence highlights a key gap: existing methods either sacrifice generality for performance or favor flexibility at the cost of accuracy.
In this work, we aim to bridge this gap by designing a quantization framework that is both tailored to the characteristics of diffusion models and broadly applicable across different architectures. Our proposed method, SegQuant, builds upon the structural insights of diffusion models but is designed with generalization and real-world deployment in mind. It combines task-specific precision with modular design, enabling seamless integration into standard toolchains while maintaining high generation quality.
Figure 1: SegQuant framework follows a top-down workflow that effectively integrates existing quantization techniques with our novel contributions.
To better understand the unique challenges of quantizing diffusion models, we begin our investigation with DiT-based architectures (Peebles and Xie 2023), which are widely adopted and representative of modern diffusion backbones. By examining the spatial semantic heterogeneity in special layers such as AdaNorm, we identify a phenomenon of semantic segmentation emerging in certain computational patterns. Additionally, we observe that the negative activation behavior introduced by activation functions like SiLU significantly impacts the final visual quality. To address these issues, we propose:
  1. 1.
    SegQuant, a unified analysis and top-down framework illustrated in Figure 1, which integrates various quantization techniques through adaptive search, extending beyond diffusion models to enhance practical versatility.
  2. 2.
    SegLinear, an adaptive segment-wise quantization method that leverages spatial and semantic variations in DiT-like architectures and generalizes effectively to other model types.
  3. 3.
    DualScale, a simple yet effective enhancement that preserves negative activations, improving generation quality while remaining compatible with standard quantization workflows.
Together, SegQuant achieves an improved trade-off between quantization accuracy and deployment flexibility. Extensive experiments demonstrate that our framework not only enhances the performance of quantized diffusion models but also generalizes well to other architectures.
Related Work
Quantization
Quantization has become a widely adopted technique for compressing deep models, with a majority of existing methods focusing on linear layers. These approaches can be broadly categorized into two types: weight-specific optimization and joint quantization of weights and activations. Weight-specific quantization techniques aim to minimize quantization error through mathematical optimization or blockwise reconstruction strategies (Nagel et al. 2020; Li et al. 2021; Frantar et al. 2022). In contrast, joint quantization methods consider both weights and activations during calibration, often leveraging smoothing or decomposition strategies to suppress outlier values (Xiao et al. 2023; Li* et al. 2025). Meanwhile, industrial quantization toolkits such as TensorRT (NVIDIA Corporation 2023) provide deployment-ready solutions, incorporating kernel-level optimizations and hardware-aware scheduling. While effective in many applications, these general-purpose toolkits often fall short when applied to inherently complex models like diffusion architectures, where domain-specific characteristics must be explicitly considered.
Quantization for Diffusion Models
The unique multi-step generative process and architectural complexity of diffusion models pose distinct challenges for quantization. Early efforts focused on timestep-dependent activation variance, proposing improved calibration schemes tailored to the denoising dynamics (Shang et al. 2023; Li et al. 2023). From an architectural perspective, several works have explored UNet-based diffusion models, designing quantization strategies that account for structural patterns such as concatenation, scaling, and temporal blocks (Li et al. 2023; Chu et al. 2024; Huang et al. 2024). More recently, attention has shifted to DiT-based diffusion models (Peebles and Xie 2023). For instance, PTQ4DiT (Wu et al. 2024) introduces channel-wise optimization through adaptive channel selection, while VIDIT-Q (Zhao et al. 2024) designs token-wise and timestep-wise techniques tailored to video generation. Despite these advances, most existing methods rely heavily on architecture-specific heuristics, limiting their generalizability across different diffusion variants and broader model families.
Our Work
In contrast to prior work, our framework does not depend on rigid architectural assumptions or diffusion-specific dynamics. Instead, SegQuant leverages intrinsic semantic structures within models, e.g., spatial heterogeneity and activation asymmetry, to improve quantization fidelity in a more principled and generalizable manner. This enables high-quality quantization while preserving compatibility with diverse model architectures and deployment pipelines.
Preliminaries
Quantization
Quantization reduces model size and computational cost by replacing full-precision floats with low-precision integers or floats. Two primary paradigms exist: quantization-aware training (QAT), which incorporates quantization effects during training, and post-training quantization (PTQ), which applies quantization after training without requiring fine-tuning. PTQ is particularly favored in real-world deployments for its simplicity and practicality.
For any given full-precision number $x$, PTQ can apply either symmetric or asymmetric quantization, both described by:
$\hat{x}=\mathrm{clip}\left(\mathrm{round}\left(\frac{x-z}{s}\right),q_{\min},q_{\max}\right),$
where $\mathrm{round}(\cdot)$ denotes rounding the input to the nearest integer; $s$, $z$, and $[q_{\min},q_{\max}]$ represent the scale, zero-point, and quantization range, respectively. The original value $x$ can be approximately recovered by the inverse operation: $x\approx s\cdot(\hat{x}+z)$.
In symmetric quantization, the zero-point $z$ is zero, simplifying the formula and hardware implementation. Asymmetric quantization allows a nonzero $z$ for greater flexibility but at the cost of increased complexity.
Diffusion Models
Diffusion models (Ho, Jain, and Abbeel 2020), such as Denoising Diffusion Probabilistic Models (DDPMs), generate high-quality samples through an iterative denoising process. The forward diffusion process gradually corrupts the input data $\mathbf{x}_{0}\sim q(\mathbf{x}_{0})$ with Gaussian noise over $T$ steps, generating latent variables $\mathbf{x}_{1},\ldots,\mathbf{x}_{T}$ according to:
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\alpha_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I}),$
where $\alpha_{t},\beta_{t}$ are hyperparameters and $\alpha_{t}=1-\beta_{t}$.
The reverse process iteratively reconstructs the original data from pure noise by estimating and removing the added noise. This is achieved using a parameterized noise prediction function $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)$, which approximates the noise residual at each step. Denoising proceeds as:
$\mathbf{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right)+\sigma_{t}\mathbf{z},$
where $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and $\sigma_{t}$ stands for the noise scale coefficient at timestep $t$.
Figure 2: Structural overview of the DiT diffusion model, highlighting latent-related modules (left) and time-related modules (right).
Recent advancements have improved efficiency through techniques such as accumulating the denoising schedule (Song, Meng, and Ermon 2022) and leveraging vector fields (Lipman et al. 2023). Modern diffusion architectures often employ Transformer-based (Vaswani et al. 2017) backbones, exemplified by the DiT family of models (Peebles and Xie 2023), which offer superior scalability and performance compared to traditional UNet structures (Ronneberger, Fischer, and Brox 2015). DiT takes as input a latent feature map $\mathbf{z}$ obtained from a pretrained VAE (Kingma and Welling 2022), splits it into non-overlapping patches, and embeds them as tokens. It incorporates temporal information through learned positional embeddings and uses Adaptive LayerNorm (Perez et al. 2018) for conditional normalization. Cross-attention layers allow for flexible conditioning on text or other modalities. Finally, a linear layer predicts the noise $\boldsymbol{\epsilon}_{\theta}$, which is used to refine the sample in the next denoising step.
Graph
Modern deep learning frameworks such as PyTorch (Paszke et al. 2019) and TensorFlow (Abadi et al. 2016) represent neural networks as directed acyclic graphs (DAGs), where nodes correspond to operations and edges denote tensor dependencies. This graph abstraction enables various graph-based optimizations such as operator fusion, memory reuse, and scheduling used by AI compilers (Chen et al. 2018; Zhang et al. 2024). In this work, we exploit the structured nature of computation graphs to guide our quantization strategy. By identifying semantically meaningful operator regions based on patterns in the graph, we enable more effective and context-aware quantization decisions.
SegQuant Framework
SegQuant Overview
In this section, we present SegQuant, a unified top-down quantization framework designed through an in-depth analysis of both diffusion-specific and general-purpose quantization techniques (Figure 1). In SegQuant, we organize existing quantization techniques into two modular components: the Optimizer (Xiao et al. 2023; Li* et al. 2025) and the Calibrator (Frantar et al. 2022), and further enhance it with our proposed modules, i.e., SegLinear and DualScale, which will be detailed in subsequent sections. In addition, the framework incorporates efficient CUDA kernels [^1] https://github.com/NVIDIA/cutlass for deployment, enabling acceleration without compromising the integrity of the quantization logic.
By combining graph-based analysis with automated configuration, SegQuant provides a flexible and extensible foundation for quantizing diverse model architectures. Although our empirical evaluation focuses on DiT-based diffusion models, the framework is model-agnostic and can be readily applied to other model architectures.
SegLinear
Key Observations for Time-Sensitive Submodules
We observe that linear layers in DiT-based diffusion models exhibit different degrees of sensitivity to quantization, depending on whether they are time-related or latent-related. As shown in Figure 2, the diffusion model consists of time-related and latent-related submodules with distinct roles. When we apply uniform INTW8A8 quantization across all layers, we find that time-related layers suffer significantly higher error, as shown in Figure 3.
Figure 3: Frobenius norm of error $\|\boldsymbol{\Delta\epsilon_{t}}\|_{F}$ over timesteps for INTW8A8 vs. FP16 across linear layers.
These findings on linear layers’ heterogeneous quantization sensitivity inspired practical optimizations. Collaborating with industry, we applied FP8 quantization to DiT-ControlNet models—where latent layers dominate computation—achieving a 1.48× speedup with negligible quality loss ($\mathrm{PSNR}>40$, $\mathrm{SSIM}>0.98$), demonstrating the benefits of architecture-aware quantization.
Figure 4: Visualization of weights in AdaNorm within the TimeEmbedding module. The distribution reveals distinct semantic patterns.
We analyzed the quantization sensitivity of linear layers within the TimeEmbedding module. As shown in Figure 4, the weights of linear transformations within the TimeEmbedding module exhibit distinct semantic partterns, which suggest that different subspaces encode semantically unique aspects of the time signal. This heterogeneity implies that operations such as Chunk and Split can implicitly split inputs into multiple semantic branches, each requiring tailored quantization treatment. Uniform quantization fails to account for these internal structures, leading to increased error. To address this issue, we propose SegLinear (Segmented quantization for linear layers), a structure-aware quantization strategy applicable across diverse architectures.
Semantics-Aware Quantization
SegLinear is a core component of the SegQuant framework, designed to reduce quantization error in computation graphs where linear operations interact with semantic partitioning patterns. These include operations such as chunk/split (output fragmentation), and stack/concat (input aggregation), which indicate that a linear layer operates over semantically distinct input or output segments. To do this, SegLinear performs fine-grained, segment-wise quantization on linear layers based on their roles in the computation graph. Specifically, it partitions the weight matrix and corresponding activations according to observed structural semantics, and applies quantization independently within each segment. SegLinear supports two primary modes:
Figure 5: SegLinear reveals two semantic patterns in the weight matrix that guide quantization.
  1. 1.
    Output-Segmented Quantization. In the following, we denote by $\hat{\mathbf{M}}$ the quantized version of any matrix $\mathbf{M}$. When the output of a linear layer is followed by operations like chunk or split, we partition the output space and apply quantization independently to each segment. Formally, given a linear transformation $\mathbf{Y}=\mathbf{XW}$, where $\mathbf{X}\in\mathbb{R}^{m\times k}$ and $\mathbf{W}\in\mathbb{R}^{k\times n}$, we decompose the weight matrix as:
    $\mathbf{W}=[\mathbf{W}_{1},\mathbf{W}_{2},\cdots,\mathbf{W}_{N}],\quad\mathbf{W}_{i}\in\mathbb{R}^{k\times d_{i}},$
    where $d_{i}$ denotes the output dimension of the $i$-th partition, such that $\sum_{i=1}^{N}d_{i}=n$. Each sub-matrix $\mathbf{W}_{i}$ is quantized separately as $\hat{\mathbf{W}}_{i}$, and the final output is constructed by concatenating segment outputs:
    $\hat{\mathbf{Y}}=[\hat{\mathbf{X}}\hat{\mathbf{W}}_{1},\hat{\mathbf{X}}\hat{\mathbf{W}}_{2},\cdots,\hat{\mathbf{X}}\hat{\mathbf{W}}_{N}].$
  2. 2.
    Input-Segmented Quantization. When the input to a linear layer comes from operations like stack or concat, we partition the input accordingly and adjust the weight matrix to match. Suppose $\mathbf{X}=[\mathbf{X}_{1},\mathbf{X}_{2},\cdots,\mathbf{X}_{N}]$, where each $\mathbf{X}_{i}\in\mathbb{R}^{m\times d_{i}}$ and $\sum_{i=1}^{N}d_{i}=k$. Then, the weight matrix is decomposed as:
    $\mathbf{W}=\left[\mathbf{W}_{1}^{\text{T}},\mathbf{W}_{2}^{\text{T}},\cdots,\mathbf{W}_{N}^{\text{T}}\right]^{\text{T}},\quad\mathbf{W}_{i}\in\mathbb{R}^{d_{i}\times n}.$
    Each segment is quantized and multiplied separately:
    $\hat{\mathbf{Y}}=\sum_{i=1}^{N}\hat{\mathbf{X}}_{i}\hat{\mathbf{W}}_{i}.$
The segment sizes $d_{i}$ are automatically inferred from the computation graph using pattern matching over operations such as chunk, split, and concat (Figure 5). This design ensures that quantization aligns with the semantic structure of the model, rather than being applied uniformly across arbitrary tensor dimensions. Our approach generalizes prior methods such as Q-Diffusion (Li et al. 2023), which targets UNet-based architectures, as special cases. By aligning quantization boundaries with structural semantics, SegLinear reduces inter-segment interference, improves fidelity, eliminates runtime group reconstruction, and enables better hyperparameter tuning.
DualScale
Polarity Asymmetric
Modern Transformer-based diffusion models, such as DiT (Peebles and Xie 2023), Stable Diffusion 3 (Esser et al. 2024), and FLUX.1 (Black Forest Labs 2024), commonly employ polarity-asymmetric activations like SiLU and GELU. Unlike ReLU, which suppresses negative values entirely, these functions retain a dense distribution of small-magnitude negative responses that are critical for preserving fine-grained semantic structure. The output distribution of SiLU and GELU is highly skewed, with wide-ranging positive values and tightly clustered negative values (see Figure 6). This imbalance poses a significant challenge for low-bit quantization, where limited bin resolution may lead to excessive compression of the negative range.
Figure 6: Activation curves of SiLU, GELU, and ReLU. The shaded regions show how SiLU and GELU retain negative values, while ReLU suppresses them.
Layer (Module) Activation Channels Neg/Pos Ratio
AdaNorm (DiT) SiLU 1536 0.955 / 0.021
AdaNorm (Ctrl.) SiLU 1536 0.645 / 0.338
FFN (DiT) GELU 6144 0.744 / 0.256
FFN (Ctrl.) GELU 6144 0.589 / 0.400
Table 1: Polarity statistics of SiLU and GELU activations from SD3.5-ControlNet on COCO, averaged over 30 timesteps. “Neg/Pos Ratio” shows the asymmetry in activation distributions.
Figure 7: Visual impact of negative-range quantization in SD3.5 (timestep 60, COCO). (a) and (c) show full images; (b) and (d) zoom in to illustrate detail and range loss.
To quantify this asymmetry, we analyzed activation outputs channel-wise across several representative layers from SD3.5-ControlNet using the COCO dataset (averaged over 30 timesteps). As shown in Table 1, a large proportion of channels consistently exhibit negative values. For instance, over 60%-70% of channels in certain AdaNorm and FFN layers remain predominantly negative during inference. However, conventional PTQ methods typically apply a single global scale (either symmetric or asymmetric) across the entire activation range. Due to the large spread of positive values, compression of the narrower negative range ($[-0.3,0]$) is often too aggressive. Even with asymmetric quantization, which shifts the zero-point, the uniform distribution of bins leads to poor resolution for negative values. Figure 7 shows that modifying only the negative activations causes subtle but perceptible detail loss, particularly in texture and background consistency. This indicates that negative activations are not residual noise but semantically meaningful components. Although overall quality remains largely intact, this underscores preserving small negative values for high-fidelity generation.
Polarity Preserving Quantization via DualScale
To address the issue of polarity asymmetry, we propose the so-called DualScale strategy. Unlike standard methods that use a single scaling factor across the full activation range, DualScale applies distinct scales to the negative and non-negative regions, thereby preserving resolution in both regions, particularly in the narrow but semantically important negative range.
Let $x\in\mathbb{R}$ denote an activation value along the computation path from an activation function to its subsequent linear operator (referred to as the act-to-linear segment). The dual-scale quantization function $Q_{\text{dual}}(x)$ is defined as:
$Q_{\text{dual}}(x)=\begin{cases}\mathrm{round}\left(\frac{x}{s_{-}}\right),&x<0\\ \mathrm{round}\left(\frac{x}{s_{+}}\right),&x\geq 0,\end{cases}$
where $s_{-}$ and $s_{+}$ denote the step sizes for the negative and positive regions, respectively. These are computed as:
$s_{-}=\frac{|\min(x)|}{q_{\min}},\quad s_{+}=\frac{\max(x)}{q_{\max}},$
where $q_{\min}$ and $q_{\max}$ denote the quantization range.
The dual-scale quantization is applied to the activation matrix $\mathbf{X}\in\mathbb{R}^{m\times k}$ in a linear layer, where $\mathbf{W}\in\mathbb{R}^{k\times n}$ is the weight matrix. Since polarity asymmetry mainly comes from activations, we apply dual-scale quantization to $\mathbf{X}$ only, keeping $\mathbf{W}$ in standard low precision. To preserve resolution and avoid destructive rounding across polarities, we decompose $\mathbf{X}$ into its non-negative and negative parts using element-wise masks:
$\mathbf{X}_{+}=\max(\mathbf{X},0),\quad\mathbf{X}_{-}=\min(\mathbf{X},0).$
The key idea is to separately quantize and process the positive and negative channels, then linearly combine the results after matrix multiplication:
$\mathbf{Y}=\mathbf{X}\mathbf{W}$ $\approx\operatorname{DeQuant}(Q_{\text{dual}}(\mathbf{X_{+}+\mathbf{X_{-}}})Q(\mathbf{W}))$
$=\left(s_{+}\cdot\mathbf{\hat{X}_{+}}+s_{-}\cdot\mathbf{\hat{X}_{-}}\right)\cdot\left(s_{w}\cdot\mathbf{\hat{W}}\right)$
$=s_{+}s_{w}\cdot\left(\mathbf{\hat{X}_{+}}\mathbf{\hat{W}}\right)+s_{-}s_{w}\cdot\left(\mathbf{\hat{X}_{-}}\mathbf{\hat{W}}\right),$
where $\mathbf{\hat{A}}=\mathrm{round}\left(\mathbf{A}/{s_{A}}\right)$ denotes the quantized version of matrix $\mathbf{A}$ scaled by $s_{A}$.
This dual-scale quantization avoids inverse zero-point correction, enabling simple output reconstruction with fixed positive and negative scales (see Appendix). It preserves small negative values often lost in standard quantization and integrates into Transformer MLPs, AdaNorm, and diffusion embeddings without retraining or custom operations.
Experiments
Setup
Datasets and Metrics
We evaluate on COCO (Lin et al. 2015), and sample 5,000 images from MJHQ-30K (Li et al. 2024) and DCI (Urbanek et al. 2024), as suggested by recent work (Li* et al. 2025). Quantization quality is measured against FP16 outputs using FID (Heusel et al. 2017), LPIPS (Zhang et al. 2018), PSNR, SSIM (Wang et al. 2004), and Image Reward (Xu et al. 2023).
Models and Hardware
We evaluate on three models: Stable Diffusion 3.5 Medium (Esser et al. 2024) (2B), FLUX.1-dev (Black Forest Labs 2024) (12B, DiT-based), and SDXL (Podell et al. 2023) (UNet-based). All experiments are run on NVIDIA RTX 4090 and L20 GPUs.
Baselines and Settings
We compare against representative PTQ baselines: Q-Diffusion (Li et al. 2023), PTQD (He et al. 2023), PTQ4DiT (Wu et al. 2024), TAC-Diffusion (Yao et al. 2024), and SVDQuant (Li* et al. 2025). For calibration, we sample 256 images for SD3 and SDXL, 64 for 8-bit FLUX, and 32 for 4-bit. All experiments use 50 sampling steps (more details see Appendix). We report results for both SegQuant-A (AMax) and SegQuant-G (GPTQ) in our experiments.
Backbone W/A Method Quality Similarity
[…]
SD3.5-DiT 8/8(int) Q-Diffusion $169.59$ $-2.072$ $0.690$ $7.92$ $0.295$
PTQD $26.53$ $0.309$ $0.520$ $10.20$ $0.417$
PTQ4DiT $16.46$ $0.752$ $0.426$ $12.18$ $0.532$
TAC $17.78$ $0.702$ $0.440$ $11.99$ $0.520$
SegQuant-A $13.90$ $0.924$ $0.384$ $12.78$ $0.563$
SegQuant-G $12.37$ $0.859$ $0.383$ $12.83$ $0.564$
4/8(int) PTQ4DiT $62.98$ $-0.190$ $0.577$ $10.06$ $0.429$
SVDQuant $20.58$ $0.725$ $0.456$ $11.76$ $0.523$
SegQuant-G $20.22$ $0.762$ $0.453$ $11.69$ $0.521$
SDXL-UNet 8/8(int) Q-Diffusion $72.81$ $-1.618$ $0.567$ $11.97$ $0.360$
PTQ4DiT $7.70$ $0.736$ $0.191$ $19.66$ $0.691$
SegQuant-A $6.39$ $0.775$ $0.141$ $21.33$ $0.742$
SegQuant-G $6.19$ $0.764$ $0.134$ $21.60$ $0.750$
8/8(fp) Q-Diffusion $5.14$ $0.897$ $0.093$ $24.31$ $0.827$
SegQuant-A $5.12$ $0.901$ $0.093$ $24.28$ $0.827$
SegQuant-G $4.83$ $0.903$ $0.082$ $24.84$ $0.838$
FLUX-DiT 8/8(int) Q-Diffusion $9.41$ $0.732$ $0.299$ $15.87$ $0.633$
PTQ4DiT $11.91$ $0.630$ $0.325$ $15.36$ $0.611$
SegQuant-A $5.89$ $0.835$ $0.150$ $19.84$ $0.770$
SegQuant-G $5.56$ $0.822$ $0.138$ $20.32$ $0.782$
4/8(int) PTQ4DiT $30.09$ $-0.039$ $0.540$ $12.09$ $0.540$
SVDQuant $7.94$ $0.783$ $0.232$ $17.29$ $0.697$
SegQuant-G $7.78$ $0.789$ $0.225$ $17.48$ $0.702$
Table 2: Main results across different backbones and models on the MJHQ-30K dataset.
Q-Diffusion
PTQD
PTQ4DiT
TAC-Diffusion
SegQuant-A
SegQuant-G
Q-Diffusion
PTQD
PTQ4DiT
TAC-Diffusion
SegQuant-A
SegQuant-G
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: thor with drone, with neon lights in the background, Intricate, Highly detailed, Sharp focus, Digital painting, Artstation, Concept art, inspired by blade runner, ghost in the shell and cyberpunk 2077, art by rafal wechterowicz and khyzyl saleem
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: Lewis Hamilton standing infront of his Formula 1 pit garage and Formula 1 car in the style of a GTA 5 grand theft auto 5 loading screen miami
Figure 8: Partial visualization of main results on the MJHQ dataset with SD3.5 and W8A8 DiT quantization.
Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: wizard bartender serving whiskey shots to a lizard patron on the moon
Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: a park with a beautiful wooden bridge over a pond, flowering trees around the banks, beautiful color correction, 16k, Alphonse Mucha style
Figure 9: Partial visualization of ablation results on the MJHQ dataset with SD3.5 and W8A8 DiT quantization. From left to right: baseline, Seg., Dual. and Seg.+Dual.
Figure 10: Performance of quantization strategies on SD3.5 (RTX 4090). INT8 (W8A8) uses SmoothQuant; INT4 (W4A8) uses SVDQuant. Model size includes only the backbone; inference time is per-step (end-to-end).
Main Results
We evaluate SegQuant on the MJHQ dataset across various models and precision levels. As shown in Table 2, it consistently achieves better image quality and higher fidelity to the original model. Additional visualizations are presented in Fig. 8, with more results in the appendix. We also analyze efficiency in Fig. 10, showing comparable memory usage to naive quantization and a modest runtime increase from segmentation and dual-scale steps, which bring notable quality gains.
Ablation Study
SegQuant enhances quantization by integrating semantic-aware optimization and calibration. Optimization methods such as SmoothQuant (Xiao et al. 2023) and SVDQuant (Li* et al. 2025) utilize SegLinear with semantic segmentation for fine-grained tuning of the hyperparameter $\alpha$. For calibration, SegQuant adopts GPTQ (Frantar et al. 2022) or AMax, where SegLinear segments weights by output features to complement GPTQ. This design yields lower quantization error, as shown in Table 3. We further isolate the effects of SegLinear and DualScale, demonstrating their individual benefits to image quality in Table 4 and Fig. 9.
Layer Name Method F-norm
w/o Seg. w/ Seg.
DiT.0.norm1 SmoothQuant 0.7041 0.5381
DiT.11.norm1 1.0684 0.9053
DiT.0.norm1 GPTQ 0.8350 0.4441
DiT.0.norm1_context 1.5166 0.7441
DiT.11.norm1 1.1719 0.9419
DiT.11.norm1_context 3.0176 1.7637
Table 3: Frobenius norm of quantization error for W8A8 single linear layers in SD3.5, comparing SmoothQuant (tunes $\alpha$, uses AMax) and GPTQ (fixed $\alpha$ = 0.5).
Method Dataset Quality Similarity
[…]
Baseline COCO 16.19 0.835 0.624 8.44 0.298
+Seg. 16.38 0.860 0.625 8.40 0.288
+Dual. 15.10 0.880 0.623 8.37 0.300
+Seg.+Dual. 14.92 0.931 0.620 8.42 0.288
Baseline MJHQ 15.92 0.814 0.422 12.41 0.531
+Seg. 15.73 0.825 0.414 12.35 0.544
+Dual. 13.07 0.884 0.400 12.64 0.547
+Seg.+Dual. 12.38 0.912 0.389 12.68 0.560
Table 4: Ablation study on COCO and MJHQ-30K with SD3.5 and W8A8 DiT quantization. SmoothQuant with $\alpha{=}0.5$.
Conclusion
We propose SegQuant, a semantics-aware quantization framework that leverages feature segmentation and polarity-sensitive scaling. It significantly enhances image quality and perceptual similarity in diffusion models based on DiT backbones, and demonstrates strong generalization to other architectures.
Appendix A Quantization Recovery Comparison
In quantized linear layers, the forward pass is typically composed of three steps: quantizing the input and weight tensors, performing matrix multiplication in the quantized domain, and then recovering the output via dequantization. Formally, for input $\mathbf{X}$ and weight $\mathbf{W}$, where $\mathbf{X}\in\mathbb{R}^{m\times k}$ and $\mathbf{W}\in\mathbb{R}^{k\times n}$:
$\mathbf{Y}\approx\mathrm{DeQuant}(\mathrm{Quant}(\mathbf{X})\cdot\mathrm{Quant}(W)),$
where $\mathrm{Quant}(\cdot)$ and $\mathrm{DeQuant}(\cdot)$ denote quantization and dequantization, respectively.
This section compares three representative quantization strategies from the perspective of how easily the original scale can be recovered after matrix multiplication:
  • Symmetric quantization: both inputs and weights are quantized using zero-centered uniform scales without offsets. Specifically, for input $\mathbf{X}$ and weight $\mathbf{W}$, the quantization is defined as:
    $\hat{\mathbf{X}}=\mathrm{round}\left(\frac{\mathbf{X}}{s_{X}}\right),\quad\hat{\mathbf{W}}=\mathrm{round}\left(\frac{\mathbf{W}}{s_{w}}\right),$
    where $s_{X}$ and $s_{W}$ are the quantization scales for input and weight, respectively. The low-bits matrix multiplication yields:
    $\hat{\mathbf{Y}}=\hat{\mathbf{X}}\hat{\mathbf{W}},$
    and the recovered output is simply:
    $\mathbf{Y}\approx s_{X}s_{W}\cdot\hat{\mathbf{Y}},$
    which is efficient and scale-preserving due to the absence of zero-points.
  • Asymmetric quantization: unlike the symmetric case, asymmetric quantization introduces non-zero offsets (zero-points), which shifts the quantized representation. For input $\mathbf{X}$ and weight $\mathbf{W}$, the quantization process is:
    $\hat{\mathbf{X}}=\mathrm{round}\left(\frac{\mathbf{X}}{s_{X}}\right)+z_{X},\quad\hat{\mathbf{W}}=\mathrm{round}\left(\frac{\mathbf{W}}{s_{W}}\right)+z_{W},$
    where $s_{X}$, $s_{W}$ are the quantization scales, and $z_{X}$, $z_{W}$ are the zero-points for input and weight, respectively.
    The low-bits matrix multiplication gives:
    $\hat{\mathbf{Y}}=\hat{\mathbf{X}}\hat{\mathbf{W}}.$
    To recover the output in full precision, we must subtract the effects of the zero-points:
    $\mathbf{Y}$ $=\mathbf{X}\mathbf{W}$
    $\approx s_{X}\left(\hat{\mathbf{X}}-z_{X}\right)\cdot s_{W}\left(\hat{\mathbf{W}}-z_{W}\right)$
    $=s_{X}s_{W}\left(\hat{\mathbf{X}}\cdot\hat{\mathbf{W}}-z_{X}\mathbf{J}_{m\times k}\hat{\mathbf{W}}-z_{W}\hat{\mathbf{X}}\mathbf{J}_{k\times n}+kz_{X}z_{W}\mathbf{J}_{m\times n}\right)$
    $=s_{X}s_{W}\left(\hat{\mathbf{X}}\cdot\hat{\mathbf{W}}-z_{X}\cdot\mathbf{J}_{m\times 1}\cdot\operatorname{rowsum}(\hat{\mathbf{W}})-z_{W}\operatorname{rowsum}(\hat{\mathbf{X}})\cdot\mathbf{J}_{1\times n}+kz_{X}z_{W}\right),$
    where $\mathbf{J}$ is a matrix of ones.
    As shown, asymmetric quantization introduces additional terms involving zero-points, which require extra additions and broadcasted summations during the recovery of the GEMM output. Although this scheme can improve accuracy when data distributions are significantly shifted, it increases computational complexity and reduces implementation efficiency.
  • Dual-scale quantization (ours): decomposes the input into positive and negative channels before quantization, preserving directional fidelity. This allows for scale-aligned matrix multiplication and avoids zero-point corrections. The final output is recovered by linearly combining the quantized results (see the dual-scale quantization formula).
Our dual-scale quantization method achieves a favorable trade-off between accuracy and computational efficiency. By decomposing the input into positive and negative channels with separate quantization scales, it preserves directional information that symmetric quantization often loses, thereby improving precision. Compared to asymmetric quantization, it avoids costly zero-point corrections and broadcasted summations during recovery. As a result, the output dequantization is simpler and more efficient, making our method well-suited for practical quantized neural network implementations.
Appendix B More Experimental Details and Results
Extra Hyperparameter Settings and Implementation Details
In all experiments involving diffusion models, we enable classifier-free guidance and set the guidance scale to 7. Regarding quantization granularity: all 8-bit quantization uses a per-tensor scheme. In 4-bit settings, weights are quantized per-channel, and activations are quantized dynamically per-token.
For SegQuant, the SmoothQuant algorithm is applied with the $\alpha$ parameter individually selected for each linear layer. We sweep $\alpha$ in the range from 0.0 to 1.0 with a step size of 0.1, choosing the value that minimizes the mean squared error (MSE) between the quantized and full-precision layer outputs. For 4-bit quantization, SegQuant uses SVDQuant as the optimizer instead of SmoothQuant. In the DualScale scheme, we focus on polarity-sensitive activation functions, specifically SiLU, GELU, and GEGLU, to ensure asymmetric activations are well preserved.
For SVDQuant, the low-rank setting is fixed at 64. For the FLUX model, singular value decomposition (SVD) is performed using float32 precision due to implementation constraints, while all other models use float64 precision for better numerical stability. For PTQD and TAC-Diffusion, calibration is applied only to the unconditional branch of the model, with 32 images used for sampling. For TAC-Diffusion, we use $\lambda_{1}{=}0.8$, $\lambda_{2}{=}0.1$, and a threshold of 4, following the original implementation. Additional implementation details and full configuration files can be found in our released codebase.
Results on Other Datasets
Beyond the main results, we report CLIP-based metrics—CLIP Score (Hessel et al. 2021) and CLIP-IQA (Wang, Chan, and Loy 2023)—on MJHQ (Table 6), and compare SegQuant with other baselines on COCO (Table 5) and DCI (Table 7), using the openai/clip-vit-large-patch14 model.
Visual Evidence from Baseline and Ablation Studies
We provide detailed visual comparisons to highlight the effectiveness of our method. These include both comparisons with existing quantization baselines and ablation studies that isolate the impact of different components in SegQuant. As shown in Figures 11, 12, 13, 14, and 15, our method better preserves semantic structure and visual fidelity, especially under aggressive quantization settings. In particular, SegQuant-G produces more consistent fine-grained textures and semantic alignment, demonstrating the benefit of integrating semantic cues into the quantization process.
Backbone W/A Method Quality Similarity
[…]
SD3.5-DiT 8/8(int) Q-Diffusion $218.04$ $-2.197$ $0.395$ $16.02$ $0.704$ $8.21$ $0.388$
PTQD $15.62$ $0.822$ $0.434$ $16.47$ $0.454$ $10.79$ $0.523$
PTQ4DiT $14.21$ $0.912$ $0.458$ $16.37$ $0.403$ $12.69$ $0.593$
TAC $16.89$ $0.863$ $0.458$ $16.31$ $0.417$ $12.53$ $0.583$
SegQuant-A $10.90$ $1.020$ $0.467$ $16.35$ $0.362$ $13.27$ $0.618$
SegQuant-G $11.06$ $0.991$ $0.457$ $16.36$ $0.376$ $13.14$ $0.601$
4/8(int) PTQ4DiT $92.36$ $0.085$ $0.411$ $16.51$ $0.570$ $10.83$ $0.504$
SVDQuant $25.48$ $0.855$ $0.454$ $16.40$ $0.432$ $12.32$ $0.582$
SegQuant-G $26.76$ $0.843$ $0.438$ $16.45$ $0.434$ $12.40$ $0.578$
SDXL-UNet 8/8(int) Q-Diffusion $74.89$ $-1.574$ $0.396$ $16.48$ $0.537$ $13.68$ $0.517$
PTQ4DiT $7.51$ $0.639$ $0.407$ $16.70$ $0.213$ $19.17$ $0.725$
SegQuant-A $5.88$ $0.645$ $0.408$ $16.71$ $0.145$ $21.32$ $0.785$
SegQuant-G $5.72$ $0.652$ $0.408$ $16.71$ $0.138$ $21.63$ $0.793$
8/8(fp) Q-Diffusion $4.94$ $0.843$ $0.428$ $16.46$ $0.104$ $23.32$ $0.832$
SegQuant1 $4.92$ $0.844$ $0.428$ $16.46$ $0.104$ $23.31$ $0.832$
SegQuant2 $4.63$ $0.839$ $0.427$ $16.46$ $0.093$ $23.90$ $0.842$
FLUX-DiT 8/8(int) Q-Diffusion $8.37$ $0.883$ $0.446$ $16.52$ $0.302$ $15.23$ $0.624$
PTQ4DiT $10.47$ $0.716$ $0.460$ $16.45$ $0.328$ $14.73$ $0.620$
SegQuant-A $5.13$ $0.907$ $0.445$ $16.52$ $0.155$ $19.40$ $0.770$
SegQuant-G $4.94$ $0.900$ $0.444$ $16.54$ $0.143$ $19.94$ $0.784$
4/8(int) PTQ4DiT $34.92$ $0.096$ $0.417$ $16.65$ $0.561$ $11.63$ $0.537$
SVDQuant $7.11$ $0.880$ $0.436$ $16.57$ $0.242$ $16.85$ $0.693$
SegQuant-G $6.86$ $0.882$ $0.439$ $16.56$ $0.232$ $17.03$ $0.700$
Table 5: COCO
Backbone W/A Method Quality W/A Method Quality
[…]
SD3.5-DiT 8/8(int) Q-Diffusion $0.444$ $15.31$ 4/8(int) PTQ4DiT 0.428 15.86
PTQD $0.441$ $15.91$
PTQ4DiT $0.461$ $15.91$ SVDQuant 0.452 15.85
TAC $0.460$ $15.90$
SegQuant-A $0.468$ $15.86$
SegQuant-G $0.466$ $15.88$ SegQuant-G 0.452 15.89
SDXL-UNet 8/8(int) Q-Diffusion $0.400$ $15.61$ 8/8(fp) Q-Diffusion 0.417 15.71
PTQ4DiT $0.430$ $15.76$
SegQuant-A $0.433$ $15.77$ SegQuant-A $0.417$ $15.71$
SegQuant-G $0.433$ $15.78$ SegQuant-G $0.418$ $15.71$
FLUX-DiT 8/8(int) Q-Diffusion $0.444$ $15.98$ 4/8(int) PTQ4DiT 0.416 15.98
PTQ4DiT $0.461$ $15.88$
SegQuant-A $0.440$ $15.94$ SVDQuant $0.435$ $15.99$
SegQuant-G $0.440$ $15.93$ SegQuant-G $0.437$ $16.01$
Table 6: MJHQ-30K
Backbone W/A Method Quality Similarity
[…]
SD3.5-DiT 8/8(int) Q-Diffusion $161.22$ $-1.973$ $0.460$ $17.92$ $0.691$ $7.80$ $0.254$
PTQD $53.38$ $-0.454$ $0.416$ $18.08$ $0.582$ $10.00$ $0.296$
PTQ4DiT $15.71$ $0.485$ $0.430$ $18.08$ $0.445$ $12.33$ $0.492$
TAC $17.22$ $0.430$ $0.432$ $18.10$ $0.461$ $12.11$ $0.478$
SegQuant-A $11.99$ $0.639$ $0.440$ $18.07$ $0.407$ $12.74$ $0.521$
SegQuant-G $11.42$ $0.639$ $0.447$ $18.06$ $0.412$ $12.40$ $0.516$
4/8(int) PTQ4DiT $67.23$ $-0.314$ $0.403$ $18.21$ $0.566$ $10.94$ $0.386$
SVDQuant $20.03$ $0.518$ $0.430$ $18.14$ $0.446$ $12.34$ $0.495$
SegQuant-G $17.16$ $0.576$ $0.435$ $18.14$ $0.438$ $12.25$ $0.497$
SDXL-UNet 8/8(int) Q-Diffusion $64.87$ $-1.466$ $0.402$ $17.65$ $0.552$ $13.64$ $0.448$
PTQ4DiT $7.15$ $0.472$ $0.415$ $17.94$ $0.197$ $19.63$ $0.686$
SegQuant-A $5.67$ $0.487$ $0.416$ $17.94$ $0.132$ $21.69$ $0.752$
SegQuant-G $5.47$ $0.504$ $0.416$ $17.94$ $0.123$ $22.11$ $0.763$
8/8(fp) Q-Diffusion $5.06$ $0.474$ $0.406$ $17.89$ $0.102$ $23.28$ $0.780$
SegQuant-A $4.96$ $0.475$ $0.406$ $17.89$ $0.101$ $23.34$ $0.782$
SegQuant-G $4.77$ $0.474$ $0.405$ $17.89$ $0.092$ $23.82$ $0.793$
FLUX-DiT 8/8(int) Q-Diffusion $8.96$ $0.513$ $0.449$ $18.01$ $0.298$ $14.40$ $0.563$
PTQ4DiT $13.04$ $0.438$ $0.472$ $17.97$ $0.552$ $9.22$ $0.321$
SegQuant-A $5.72$ $0.592$ $0.452$ $18.01$ $0.149$ $18.29$ $0.715$
SegQuant-G $5.32$ $0.590$ $0.451$ $18.02$ $0.131$ $18.97$ $0.737$
4/8(int) PTQ4DiT $60.24$ $-0.393$ $0.418$ $17.88$ $0.592$ $10.98$ $0.465$
SVDQuant $8.12$ $0.575$ $0.442$ $18.03$ $0.224$ $16.19$ $0.635$
SegQuant-G $8.56$ $0.566$ $0.446$ $18.03$ $0.219$ $16.31$ $0.639$
Table 7: DCI
Q-Diffusion
PTQD
PTQ4DiT
TAC-Diffusion
SegQuant-A
SegQuant-G
Q-Diffusion
PTQD
PTQ4DiT
TAC-Diffusion
SegQuant-A
SegQuant-G
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: Detailed photography of Pokemon Clefable pick up on the ground some moondust. The final image should be ultradetailed. colors graded with a muted palette of fresh and warm tones to convey the dramatic, exciting essence of the scene. Enhance the captivating charadesign with postproduction techniques such as tone mapping, chromatic aberration, and ambient occlusion for a polished, evocative masterpiece, happy hour, cinematic, award winning photo of the year, dramatic lighting, photorealistic, 50x65, 8k
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: Mario in a 3D Minecraft world, blocky world, Minecraft world, minecraft trees, sunset in the horizon, colourful lighting, centered, front view, red hat, white gloves, blue pants, moustache, ultrarealistic, unreal engine 5, HDR, Ray tracing, cinematic, depth of field, sharp focus, natural light, concept art, super resolution, cartoon style, pixar style, flat coloring.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: Porcelain antique Cup of tea, multicolored cakes, samovar, a bouquet of lilies of the valley, a chocolate bar, an alarm clock, the sun, joy, a beautifully embroidered chintz tablecloth, the sun shines brightly outside the window
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: morning coffee, morning light, unreal engine
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: Mongolian nature with blue sky, mountain, steppe, Mongolian traditional yurt, hyperrealistic, photorealistic, cinematic colour grading, cinematic frame composition, 8K, shot on konica centuria film Leica M6
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: This world, a part of infinity, Contains an endless variety. Yet it stands alone, not part of any, An enigma, a mystery.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: hyperrealistic image of Harry Potter poster, lit by colored lights, in a case made in transparent plastic, full square picture, light background, extreme details, cinematic, vector, 8K resolution, HD, 34 view, high detail
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: wolf lounging in surreal scene of forest, drawn by johnen vasquez, pop art, high impact, wild and vivid colors, neon accents, blacklight reflective, magic and wonder, vivid colors, psychedelic, super art, high detail
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: thor with drone, with neon lights in the background, Intricate, Highly detailed, Sharp focus, Digital painting, Artstation, Concept art, inspired by blade runner, ghost in the shell and cyberpunk 2077, art by rafal wechterowicz and khyzyl saleem
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: hecate walking in a temple forest, anime, magical realism, hd, photorealistic, anime
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: modern one store house buried in the green mountain in forest cliff rectangle long facade with a deck, with black stone, inside of a mountain, warm light, perspective view with a reflexion in a lake
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: oil painting in sally swatland style, boy and girl holding hands, walking along the sea shore, sea gulls flying
Figure 11: More visualization of main results on the MJHQ dataset with SD3.5 and W8A8 DiT quantization.
Q-Diffusion
PTQ4DiT
SegQuant-A
SegQuant-G
Q-Diffusion
PTQ4DiT
SegQuant-A
SegQuant-G
Q-Diffusion
PTQ4DiT
SegQuant-A
SegQuant-G
Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: ukiyoe painting of tokyo city, colorful, high resolution
Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: Photorealistic, Realistic aerial view 3D big render soccer stadium, asian style, summer time, fantastical lightning, in the middle of an big asian city, full detailed, resulting in an intricately detailed and wondrous image, 8k.
Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: a luxury bedroom with a view of paris
Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: a minimalistic 3 colored and artistic illustration that describes a travel company without any letters or text included. make it in the format of 916
Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: photograph of a private Detective, age 41, irish, big, reddish hair, ruddy skin complexion, taken with a fujifilm xt5, ultra realistic, highly detailed, 8k
Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: broken plane parked on snowy mountains with burning debris, Bright color palette, 150mm, realistic, gritty, sun shafts, v5
Figure 12: More visualization of main results on the MJHQ dataset with SDXL and W8A8 DiT quantization.
Q-Diffusion
PTQ4DiT
SegQuant-A
SegQuant-G
Q-Diffusion
SVDQuant
SegQuant-G
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Prompt: a cat, its face looks like Winston Churchill, standing up like a human, in a black and white photo, looking out of a window.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Prompt: an owl sits on a rock and looks outward, in the style of nikon d850, light beige and amber, exaggerated poses, explosive wildlife, dansaekhwa, multiple points of view.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Prompt: medium sized kitchen, bombay furniture, horizontal ombre yellow to green, French provincial style, braai in central hearth, refrigerator, large butcher block, golden hour.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Prompt: a lasagne outside an italian restaurant in the city bologna at midday with bright light. Ultra High Definition. High detailed. HD. Photorealistic.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Prompt: medium sized kitchen, bombay furniture, horizontal ombre yellow to green, French provincial style, braai in central hearth, refrigerator, large butcher block, golden hour.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Prompt: birdseye view of Mediterranean castle with Romanesque revival influences, photorealistic, 8k highest resolution, horizon line visible in the back, rocky cliff tropical and clear sky.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Prompt: majestic islamic mosque in the mountain landscape beautiful.
Figure 13: Additional visualizations of the main results on the MJHQ dataset, comparing FLUX and W8A8 DiT (left) with W4A8 (right) quantization.
FP16
Q-Diffusion(8/8)
PTQ4DiT(8/8)
SegQuant-A(8/8)
SegQuant-G(8/8)
Q-Diffusion(4/8)
SVDQuant(4/8)
SegQuant-G(4/8)
Refer to caption Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Prompt: A body of water it is brown and green the sky is bright and there are many clouds in it A bright day the sky is very blue also there are clouds here that are very thick and very full also in the center of the sky is the biggest clouds. In the center of the image there is a large body of water that is green and blue. Going across the body of water is a large gray bridge with four towers on it two on each side and connecting each set of towers is a large beam that is a walkway. The walkway has windows that allow looking out over the bay. There is a large bed of rocks that is close to the camera of different colors dark gray brown white and light gray also there is also a yellow water bottle that has ben deposited amongst the rocks.
Refer to caption Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Prompt: This is a small room with blue walls inside of a church with a white alter with elaborate gold trim in the center. Two gold stands are in the room near the front of the alter. There is a bay window in the back. There is a plant in front of each window. The top of the alter is round with gold trim and has a gold cross on the top. A light is hanging on the top right side of the room.
Refer to caption Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Prompt: A brick road that starts in the foreground goes down a slope and into the background. There are buildings on the left and right sides of the road. A brick road that starts in the foreground extends back and to the left into the background. There is a white vehicle traveling along the road. Buildings can be seen lining the area to the right of the road. There is a field to the left of the road, and buildings to the left of it in the background. There is a religious building in the background that has two towers rising up from it. A wooded area can be seen on the right hand side of the background. A mountainside can be seen on the left hand side of the background. There are trees growing on the mountainside, and mist in the air to the right of the mountainside.
Refer to caption Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Prompt: Tourist merchandise is hanging on a wall on display for people to buy. There are shirts on the left, magnets in the middle, and various objects on the right. Five rows of various shirts that are folded into squares and inside of clear plastic are on the left side of the image. A wall in the middle section is covered in rows of bright colored magnets for sale. There is a round yellow and black price tag in the magnet section. The right side of the image has shelves at the top with miniature figures of the clock tower and red phone booth. Four rows of keychains are hanging below the miniature figures. A shelf with stacks of toy taxis and busses are below the key chains. A book is on its side on a shelf below the toys along with round silver plates. A shelf filled with various sizes of snow globes is below the book. The bottom shelf on the right has decorative plates on stands. A line of mugs is between the magnets and the items on the right.
Refer to caption Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Prompt: This is a roll of a fabric that is multiple shades of blue and has a leaf pattern woven into it. It is surrounded by other bright and intricate patterned rolls of fabric. This is a roll of a light blue with dark blue pattern, woven fabric surrounded by rolls and stacks of folded brightly patterned fabrics.
Refer to caption Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Prompt: A rectangular sign that contains the logo for Aldi grocery stores can be seen on the side of a building. There are several trees rising up behind the building. A blue-colored wall of a building that has several lines going down its surface can be seen going across the bottom half of the image. There is a rectangular, white-colored sign hanging off of the wall. The sign has a red border going around its edges. There is a large letter “A” in the middle of the sign that is made out of curving blue and light blue stripes. There is a blue rectangle going across the bottom of the sign that has “ALDI” going across it in white lettering. There is a white-colored object on the right side of the wall that has a rounded red top. Several trees can be seen rising up from behind the building on the right hand side of the image. It is daytime, and the sky above has gray clouds floating across its surface.
Figure 14: Additional visualizations of the main results on the DCI dataset, comparing FLUX and W8A8 DiT with W4A8 quantization.
Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: cartoon, vintage poster look, poke bowl, salmon, with fresh vegetables, with a bowl of rice, hand drawn, tropical sunset, chopsticks, 4k
Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: Star Wars poster design by Obey, retro 1980 street color propaganda style, arrangement with golden ratio, high quality render texture
Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: Castle made of smoke, castle, Evil, scary, dark, ghostly, green aurora borealis in the sky, spooky, intricate detail, high resolution, atmospheric, 4K HDR, Film Still, Hyperrealistic, Disney movie
Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: an old black mans face painted on a wall with some cracks, in the style of topographical realism, grit and grain, dorothea tanning, matte photo, stone, symbolic images, grungy patchwork, wide angle view
Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: kitchen design thinking studio, cinematic lighting, phorealistic, 8k
Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: lotus blossoms ghibli style animation
Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: shibainu, dog, japan, dog on the bridge, behind of bridge exists japanese temple, leaves changing color, ziburi style
Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: temples of various religions stand as separate buildings in a beautiful forest on which sunlight falls and birds fly
Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: Create a hyperrealistic portrayal of arabesque architecture, showcasing the elaborate archways and domes adorned with intricate carvings and mosaics. Ensure that the camera lens captures the depth and shadows of the design, highlighting the architectural details.
Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: pretty anime girl. Kawaii Face Style. flowing black hair. sparkling blue eyes. dancing in a dreamy field of flowers. rays of shimmering light. polar stratospheric clouds in the background. wearing bohemianchic clothing. braids and flowers in her hair. vibrant hues. distinctive and eyecatching fine art. splash watercolor. masterpiece
Refer to caption Refer to caption Refer to caption Refer to caption
Prompt: beautiful nature scene with dew dripping from flowers The photo was skillfully taken with a Nikon camera. The D850 DSLR paired with the versatile Nikkor 2470mm f2.8 lens, renowned for its sharpness and exceptional color reproduction. The f8 aperture is chosen to provide a deep depth of field and sharp detail capture of the entire scene. The ISO sensitivity is set to 200 and the shutter speed is 1500 second. photography uses bright, natural sunlight reflecting off a lake, illuminating the entire scene with harsh, cool light and highlighting the contrasting shadows that define the contours of the landscape.
Figure 15: More visualization of ablation results on the MJHQ dataset with SD3.5 and W8A8 DiT quantization. From left to right: baseline, Seg., Dual. and Seg.+Dual.